OpenStudy (nuccioreggie):

A window is 4 feet wide and 8 feet tall. A rectangle is shown. The length of the rectangle is labeled 4 feet. The width of the rectangle is labeled 8 feet. A manufacturer wants to use a scale factor of 1.5 to enlarge a window. What will the area of the window be after it is enlarged? 32 ft2 48 ft2 52.25 ft2 72 ft2

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (hockeychick23):

(1.5*8)+(1.5*4)= ?

OpenStudy (anonymous):

(4*1.5)(8*1.5)

OpenStudy (nuccioreggie):

I think its C

OpenStudy (anonymous):

no.....4*1.5=6 8*1.5=12 6*12=72

OpenStudy (nuccioreggie):

So the correct answer is D